For nitrogen the molar volume at s.t.p is 0.0024m³ and the relative molecular mass is 28. Calculate the root mean square molecular speed at s.t.p
1
Expert's answer
2020-11-11T08:56:50-0500

"V_m=\frac{RT}{p}"

"v=\sqrt{\frac{3RT}{M}}=\sqrt{\frac{3pV_m}{M}}=\sqrt{\frac{3\cdot1.01\cdot10^5\cdot2.4\cdot10^{-3}}{28\cdot10^{-3}}}=161" "\frac ms"